OpenStudy (anonymous):

need help with systems of equations 2x-9y=32 4x+y=-12 (I need to find x and y)

OpenStudy (yanasidlinskiy):

Do you know how to solve for either one?

OpenStudy (anonymous):

not really :(

OpenStudy (yanasidlinskiy):

I guess texaschic101 has gotten it....I'll let them have it.

OpenStudy (anonymous):

lol thank you anyway :)

OpenStudy (texaschic101):

2x - 9y = 32 4x + y = -12 ---> y = -4x - 12 now sub -4x - 12 in for y in the first equation 2x - 9(-4x - 12) = 32 -- distribute 2x + 36x + 108 = 32 -- subtract 108 from each side 2x + 36x = 32 - 108 -- combine like terms 38x = - 76 -- divide both sides by 38 x = - 2 now sub -2 in for x in either equation to find y 4x + y = -12 4(-2) + y = -12 -8 + y = -12 y = -12 + 8 y = - 4 check.. 2x - 9y = 32 2(-2) - 9(-4) = 32 -4 + 36 = 32 32 = 32 (correct) x = -2 and y = -4

OpenStudy (anonymous):

THANK YOU so muuch :)!!!